OpenStudy (anonymous):

5x-2y=14 -x-3y=4 ...solve by the method of elimination

OpenStudy (anonymous):

5x -2y =14 -5x-15y =20 -17y =34 y = -2 Plug y into the second equation and solve for x.

OpenStudy (anonymous):

how do i plug -2 onto the second equation, can you teach me

OpenStudy (callisto):

Put y=-2 into -x-3y=4 -x -3(-2) = 4 -x + 6 = 4 Solve x

OpenStudy (anonymous):

-x=-2

OpenStudy (anonymous):

?

OpenStudy (anonymous):